OpenStudy (anonymous):

integral of 1/sqrt(2x+1)

OpenStudy (anonymous):

this is equal to:\[\int\limits_{}^{}(2x - 1)^{-1/2} dx\]

OpenStudy (anonymous):

yes but how ?

OpenStudy (anonymous):

let u = 2x + 1 du = 2 dx \[\frac{ 1 }{ 2 } \int\limits_{}^{} u^{-1/2}du\] integrate and plug u back in

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

glad i could help :)